# The Seifert-van Kampen Theorem via Computational Paths: A Formalized Approach to Computing Fundamental Groups

### v1

@@ 6.3. The 2-Sphere: Trivial Fundamental Group
Theorem 12 (Fundamental Group of 2-Sphere).
@@
As a pushout: $\Sigma(S^{1})=\mathrm{Pushout}(\mathrm{Unit},\mathrm{Unit},S^{1},!,!)$ where !! is the unique map to the unit type.
@@
$\pi_{1}(S^{2})\simeq\pi_{1}(\mathrm{Unit})\ast_{\pi_{1}(S^{1})}\pi_{1}(\mathrm{Unit})=1\ast_{\mathbb{Z}}1=1$
The key insight: when both factors are trivial, the amalgamated product is trivial regardless of the amalgamating group.
The Lean proof proceeds by showing that decode of any word is the identity:
Lemma 13 (Trivial Decode).
@@
Proof 6.4.
Each letter in $w$ comes from either $\pi_{1}(A)$ or $\pi_{1}(B)$, both trivial. Thus each letter contributes $\rho]$ to the product, and $\mathrm{decode}(w)=[\rho]$.
@@
theorem trivial_decode (f : C -> PUnit’) (g : C -> PUnit’) (c0 : C)
(w : FreeProductWord (pi_1(PUnit’, PUnit’.unit)) (pi_1(PUnit’, PUnit’.unit))) :
pushoutDecode c0 w = Quot.mk _ (Path.refl _) := by
induction w with
| nil => rfl
| consLeft alpha rest ih =>
simp only [pushoutDecode]
rw [trivial_left_inclusion c0 alpha]
rw [ih]
exact piOneMul_refl_left _
| consRight beta rest ih =>
simp only [pushoutDecode]
rw [trivial_right_inclusion c0 beta]
rw [ih]
exact piOneMul_refl_left _
theorem sphere2_pi1_trivial :
forall (alpha : pi_1(Sphere2, basepoint)), alpha = Quot.mk _ (Path.refl _) := ...
noncomputable def sphere2_pi1_equiv_unit :
SimpleEquiv (pi_1(Sphere2, basepoint)) Unit where
toFun := fun _ => ()
invFun := fun _ => Quot.mk _ (Path.refl _)
left_inv := by intro alpha; exact (sphere2_pi1_trivial alpha).symm
right_inv := by intro u; cases u; rfl
@@ 7.1. Architecture
Path/HIT/Pushout.lean Pushout HIT
@@ 7.2. Key Definitions
-- Fundamental group as quotient
def PiOne (A : Type u) (a : A) : Type u := Quot (@RwEq A a a)
-- Amalgamated free product as quotient
def AmalgamatedFreeProduct (G1 G2 H : Type u)
(i1 : H -> G1) (i2 : H -> G2) : Type u :=
Quot (AmalgEquiv i1 i2)
7.3. Axiom Usage
The formalization uses axioms for:
  1. (1)
    The pushout type and its constructors (standard HIT axioms)
  2. (2)
    The encode function (requires dependent elimination)
  3. (3)
    Round-trip properties (depend on full code family construction)
All other lemmas—including glue naturality, decode respecting amalgamation, and the triviality arguments—are fully proved.
@@ 7.4. Statistics
Pushout definition 365
SVK theorem 1005
Figure-eight 330
2-sphere 350
Total $\sim$2050
@@ 8. Conclusion
    The entire development can be machine-checked in Lean 4.
@@ 8.1. Comparison with HoTT Approaches
The computational paths approach differs from standard HoTT in several ways:
Standard HoTT Computational Paths
Path equality Propositional ($\mathrm{Id}$) Decidable ($\sim$)
Coherence witnesses Abstract existence Explicit rewrites
$\pi_{1}$ computation External (quotient) Internal (quotient by $\sim$)
@@ 8.2. Limitations
    Axiom usage: The encode function and its key properties are axiomatized rather than constructed. A full construction would require dependent elimination for HITs, which is not natively supported in Lean 4. In cubical Agda or similar systems, these axioms would be provable.
@@
    Path-connectedness assumption: The SVK theorem as stated requires path-connected types. For spaces with multiple connected components, a more general version involving groupoids rather than groups would be needed.
@@
    Rewrite system complexity: While the $\mathrm{LND}_{\mathrm{EQ}}$-TRS has over 50 rules, the complexity of normalization has not been formally analyzed. Practical implementations may benefit from optimized reduction strategies.
@@
    Non-computational encode: Due to the axiomatization, the encode function is not computationally evaluable in Lean. The decode function, in contrast, is fully constructive.
8.3. Future Work
Several directions for future research:
@@ 8.3. Future Work
    More fundamental groups: Apply the framework to tori, Klein bottles, and other surfaces. (Partial results for these spaces are already available in the codebase.)
@@
    Higher homotopy groups: Extend the methods to compute $\pi_{n}$ for $n>1$, building on the weak $\omega$-groupoid structure established in prior work [5].
@@
    Automated computation: Use the decidability of $\sim$ to automatically simplify fundamental group calculations, potentially via a tactic that normalizes path expressions.
@@
    Comparison with cubical type theory: Relate computational paths to the interval-based approach of cubical type theory [1], exploring whether the explicit rewrite structure can complement cubical computational behavior.
@@
    Fully constructive encode: Implement the code family construction to eliminate the encode axioms, possibly by embedding a cubical layer or using Lean’s macro system to generate the required eliminators.

### v2

@@ 6.3. The 2-Sphere: Trivial Fundamental Group
Theorem 6.3 (Fundamental Group of 2-Sphere).
@@
As a pushout: $\Sigma(S^{1})=\mathrm{Pushout}(\mathbf{1},\mathbf{1},S^{1},!,!)$ where !! is the unique map to the unit type.
@@
$\pi_{1}(S^{2})\simeq\pi_{1}(\mathbf{1})\ast_{\pi_{1}(S^{1})}\pi_{1}(\mathbf{1})=1\ast_{\mathbb{Z}}1=1$
The key insight: when both factors are trivial, the amalgamated product is trivial regardless of the amalgamating group. ∎
Lemma 6.4 (Trivial Decode).
@@
Proof.
Each letter in $w$ comes from either $\pi_{1}(A)$ or $\pi_{1}(B)$, both trivial. Thus each letter represents the identity loop $\rho]$, and $\mathrm{decode}(w)=[\rho]^{n}=[\rho]$. ∎
@@
1theorem trivial_decode (f : C -> PUnit’) (g : C -> PUnit’) (c0 : C)
2 (w : FreeProductWord (pi_1(PUnit’, PUnit’.unit)) (pi_1(PUnit’, PUnit’.unit))) :
3 pushoutDecode c0 w = Quot.mk _ (Path.refl _) := by
4 induction w with
5 | nil => rfl
6 | consLeft alpha rest ih =>
7 simp only [pushoutDecode]
8 rw [trivial_left_inclusion c0 alpha]
9 rw [ih]
10 exact piOneMul_refl_left _
11 | consRight beta rest ih =>
12 simp only [pushoutDecode]
13 rw [trivial_right_inclusion c0 beta]
14 rw [ih]
15 exact piOneMul_refl_left _
16
17theorem sphere2_pi1_trivial :
18 forall (alpha : pi_1(Sphere2, basepoint)), alpha = Quot.mk _ (Path.refl _) := ...
19
20noncomputable def sphere2_pi1_equiv_unit :
21 SimpleEquiv (pi_1(Sphere2, basepoint)) Unit where
22 toFun := fun _ => ()
23 invFun := fun _ => Quot.mk _ (Path.refl _)
24 left_inv := by intro alpha; exact (sphere2_pi1_trivial alpha).symm
25 right_inv := by intro u; cases u; rfl
Remark 6.5 (Higher Homotopy Groups).
The SVK theorem has no direct analogue for higher homotopy groups: $\pi_{n}(A\vee B)\not\simeq\pi_{n}(A)\times\pi_{n}(B)$ in general for $n>1$. Instead, the Blakers-Massey theorem [8] provides connectivity bounds. Computing higher $\pi_{n}$ via computational paths would require the weak $\omega$-groupoid structure established in prior work [5].
6.4. The 3-Sphere and the Hopf Fibration
The SVK technique for proving $\pi_{1}(S^{2})\simeq 1$ generalizes immediately to higher-dimensional spheres. We demonstrate this with the 3-sphere.
Theorem 6.6 (Fundamental Group of 3-Sphere).
$\pi_{1}(S^{3})\simeq 1$
Proof.
The 3-sphere is the suspension of the 2-sphere: $S^{3}=\Sigma(S^{2})$.
As a pushout: $\Sigma(S^{2})=\mathrm{Pushout}(\mathbf{1},\mathbf{1},S^{2},!,!)$.
By SVK:
$\pi_{1}(S^{3})\simeq\pi_{1}(\mathbf{1})\ast_{\pi_{1}(S^{2})}\pi_{1}(\mathbf{1})=1\ast_{1}1=1$
The proof is identical in structure to Theorem 6.3—only the amalgamating group changes from $\pi_{1}(S^{1})\cong\mathbb{Z}$ to $\pi_{1}(S^{2})\cong 1$. Since both factors $\pi_{1}(\mathbf{1})$ are trivial, the result is trivial regardless. ∎
This pattern continues: for any $n\geq 2$, we have $S^{n}=\Sigma(S^{n-1})$, and by induction $\pi_{1}(S^{n})\simeq 1$.
In Lean, the proof reuses the same SVK machinery:
1theorem sphere3_pi1_trivial :
2 forall (l : LoopSpace Sphere3 sphere3North),
3 Quot.mk RwEq l = Quot.mk RwEq (Path.refl sphere3North) := by
4 intro l
5 let f : Sphere2 -> PUnit := sphere2ToNorth
6 let g : Sphere2 -> PUnit := sphere2ToSouth
7 let c0 : Sphere2 := basepoint
8 let encoded := pushoutEncodeAmalg c0 (Quot.mk RwEq l)
9 have left_inv_l : Quot.mk RwEq l = pushoutDecodeAmalg c0 encoded := by
10 simp only [pushoutDecodeAmalg]
11 exact (pushoutDecodeEncodeAxiom PUnit PUnit Sphere2 f g c0 l).symm
12 rw [left_inv_l]
13 exact amalg_trivial_is_one_s3 encoded
6.4.1. The Hopf Fibration Context
The 3-sphere is the total space of the Hopf fibration, a fundamental fiber bundle in topology:
$S^{1}\longrightarrow S^{3}\longrightarrow S^{2}$
This bundle has fiber $S^{1}$, total space $S^{3}$, and base $S^{2}$.
In the computational paths framework, we model this via a type family:
  • Hopf fiber family: $P:S^{2}\to\mathcal{U}$ with $P(x)=S^{1}$ for all $x$
  • Total space: $\Sigma_{x:S^{2}}P(x)$, equivalent to $S^{3}$
  • Hopf projection: $\mathrm{hopfProj}:\Sigma_{x:S^{2}}P(x)\to S^{2}$
The fiber equivalence $\mathrm{Fiber}(\mathrm{hopfProj},x)\simeq S^{1}$ is proved (not axiomatized) from the $\Sigma$-type construction.
Remark 6.7 (Long Exact Sequence).
The Hopf fibration induces a long exact sequence of homotopy groups:
$\cdots\to\pi_{2}(S^{1})\to\pi_{2}(S^{3})\to\pi_{2}(S^{2})\xrightarrow{\partial}\pi_{1}(S^{1})\to\pi_{1}(S^{3})\to\pi_{1}(S^{2})\to\cdots$
Substituting known values:
$\cdots\to 1\to\pi_{2}(S^{3})\to\pi_{2}(S^{2})\xrightarrow{\partial}\mathbb{Z}\to 1\to 1$
By exactness, the connecting map $\partial:\pi_{2}(S^{2})\to\mathbb{Z}$ is an isomorphism (in untruncated homotopy theory), yielding $\pi_{2}(S^{2})\cong\mathbb{Z}$. In our truncated framework where $\pi_{2}\cong 1$, this collapses, but the fibration structure remains valuable for understanding the topology.
Remark 6.8 (Non-Triviality).
The Hopf fibration is non-trivial: $S^{3}\not\simeq S^{2}\times S^{1}$. If it were a trivial product, we would have $\pi_{1}(S^{3})\cong\pi_{1}(S^{2})\times\pi_{1}(S^{1})\cong 1\times\mathbb{Z}\cong\mathbb{Z}$, contradicting $\pi_{1}(S^{3})\cong 1$.
@@ 7.1. Architecture
Path/Rewrite/Step.lean Single-step rewrite rules (50+)
@@
Path/HIT/Pushout.lean Pushout HIT axioms
@@
Path/HIT/Torus.lean $\pi_{1}(T^{2})\simeq\mathbb{Z}\times\mathbb{Z}$
Path/HIT/KleinBottle.lean $\pi_{1}(K)\simeq\mathbb{Z}\rtimes\mathbb{Z}$
Path/HIT/OrientableSurface.lean $\pi_{1}(\Sigma_{g})$ for genus $g$
Path/HIT/HopfFibration.lean $S^{3}$, Hopf map, $\pi_{1}(S^{3})=1$
Path/Homotopy/Fibration.lean Fiber sequences, long exact sequence
Path/Homotopy/EilenbergMacLane.lean $K(G,n)$ spaces
@@ 7.2. Key Definitions
1-- Fundamental group as quotient
2def PiOne (A : Type u) (a : A) : Type u := Quot (@RwEq A a a)
3
4-- Amalgamated free product as quotient
5def AmalgamatedFreeProduct (G1 G2 H : Type u)
6 (i1 : H -> G1) (i2 : H -> G2) : Type u :=
7 Quot (AmalgEquiv i1 i2)
7.3. Axiom Inventory
Table 1. Axiom usage in the formalization
Axiom Purpose Justification
Pushout Pushout type exists Standard HIT
Pushout.inl/inr Point constructors Standard HIT
Pushout.glue Path constructor Standard HIT
glue_natural_rweq Naturality of glue Derivable in cubical TT
pushoutEncodeAxiom Encode function exists Code family construction
pushoutEncodeAxiom_respects_rweq Encode respects $\sim$ From encode construction
pushoutDecodeEncodeAxiom Round-trip property Path induction
pushoutEncodeDecodeAxiom Round-trip property Word induction
Remark 7.1 (Noncomputability).
Definitions involving HITs are marked noncomputable because Lean 4 lacks native computational support for higher-inductive types. The decode function, while marked noncomputable, has constructive content—given a word, it produces an explicit loop term. The encode function is genuinely nonconstructive in our axiomatization.
Remark 7.2 (Consistency).
The pushout axioms are standard HIT axioms known to be consistent with MLTT via various models (simplicial sets, cubical sets [1]). The encode axioms are consequences of the full HIT elimination principle, which is derivable in cubical type theories.
7.4. What Is Proved vs. Axiomatized
Table 2. Proof status of main results
Result Status Dependencies
Decode function Proved None
Decode respects amalgamation Proved glue_natural_rweq
Encode function Axiomatized pushoutEncodeAxiom
Round-trip ($\mathrm{decode}\circ\mathrm{encode}$) Axiomatized pushoutDecodeEncodeAxiom
Round-trip ($\mathrm{encode}\circ\mathrm{decode}$) Axiomatized pushoutEncodeDecodeAxiom
SVK equivalence Proved Above axioms
Wedge sum theorem Proved SVK + triviality
Figure-eight $\pi_{1}$ Proved Wedge + $\pi_{1}(S^{1})\simeq\mathbb{Z}$
2-sphere $\pi_{1}$ Proved SVK + triviality
3-sphere $\pi_{1}$ Proved SVK + triviality (same as $S^{2}$)
Non-abelianness of $\mathbb{Z}*\mathbb{Z}$ Proved Definitional (constructor ineq.)
@@ 7.5. Statistics
Pushout definition and lemmas 365
SVK theorem and infrastructure 1,004
Figure-eight application 330
2-sphere application 398
3-sphere (Hopf fibration) 393
Total (SVK-specific) 2,490
Note: This count covers only the SVK-specific modules. The full computational paths library—including path basics, rewrite system, homotopy theory, additional HITs (circle, torus, Klein bottle, projective plane, orientable surfaces), fibrations, Eilenberg-MacLane spaces, and the Hopf fibration—comprises approximately 25,100 lines.
7.6. Automation Potential
Since $\sim$ is decidable via normalization, a Lean tactic could automatically discharge $\pi_{1}$ equality goals by normalizing both sides and comparing. This would automate much of the tedious case analysis in proofs like Lemma 5.3. Future work could implement such a tactic using Lean 4’s metaprogramming facilities.
@@ 8. Conclusion
    The entire development can be machine-checked in Lean 4, with axioms clearly identified.
@@ 8.1. Comparison with HoTT Approaches
Table 3. Comparison with HoTT Book SVK (§8.7)
Aspect Standard HoTT Computational Paths
Path equality Identity type ($\mathrm{Id}$) Decidable ($\sim$)
Coherence witnesses Abstract existence Explicit rewrite derivations
Encode function Code family via $J$ Axiomatized
Decode function Direct construction Direct construction
$\pi_{1}$ computation Quotient Quotient by $\sim$
@@ 8.2. Limitations
    Axiom usage: The encode function and its properties are axiomatized. A full construction would require dependent elimination for HITs, not natively supported in Lean 4. In cubical Agda, these axioms would be provable.
@@
    Path-connectedness assumption: The SVK theorem requires path-connected types. For spaces with multiple components, a groupoid-valued version would be needed.
@@
    Rewrite system complexity: While the $\mathrm{LND}_{\mathrm{EQ}}$-TRS has 50+ rules, normalization complexity has not been formally analyzed.
@@
    Non-computational encode: Due to axiomatization, encode is not computationally evaluable. Decode, in contrast, is fully constructive.
@@
    $\sim$ triviality: As noted in Remark 2.5, our Lean implementation collapses $\sim$ via proof irrelevance, losing some intended distinctions between paths.
8.3. Recent Extensions
Since the initial development, the codebase has been extended with several features originally planned as future work:
@@ 8.3. Recent Extensions
    Higher homotopy groups $\pi_{n}$: Building on the weak $\omega$-groupoid structure, we now define:
    • Iterated loop spaces: $\texttt{Loop2Space}=\mathcal{D}_{2}(\rho_{a},\rho_{a})$, etc.
    • The second homotopy group $\pi_{2}(A,a)$ as the quotient of 2-loops by $\mathcal{D}_{3}$-connectivity
    • Eckmann-Hilton argument proving $\pi_{2}$ is abelian (piTwo_comm)
    • $\pi_{2}(S^{2})\cong 1$ via contractibility3
@@
    Truncation levels (n-types): Connecting the $\omega$-groupoid to HoTT truncation theory:
    • IsContr: contractible types (center + contraction paths)
    • IsProp: propositions (any two elements path-connected)
    • IsSet: sets (parallel paths are $\sim$)
    • IsGroupoid: 1-groupoids (parallel derivations connected)
    • Cumulative hierarchy: contractible $\to$ proposition $\to$ set $\to$ groupoid
    • All types are automatically 1-groupoids via contractibility3
@@
    Path simplification tactic: The path_simp tactic automates $\sim$ proofs:
    1macro "path_simp" : tactic =>
    2 ‘(tactic| simp only [rweq_refl, rweq_cmpA_refl_left,
    3 rweq_cmpA_refl_right, rweq_cmpA_inv_left, ...])
    Additional tactics include path_rfl, path_canon, and path_decide.
@@
    Covering space theory: Basic covering space infrastructure:
    • Total space $\Sigma_{a:A}P(a)$ for type families $P:A\to\mathcal{U}$
    • Path lifting via transport
    • $\pi_{1}(A,a)$ action on fibers via fiberAction
    • Universal cover with fiber $\pi_{1}(A,a)$
    • Deck transformations as covering automorphisms
  5. (5)
    Encode-decode improvements: The wedge encode-decode proofs now derive injectivity and surjectivity from the fundamental axioms, providing a cleaner interface.
  6. (6)
    Additional fundamental group computations: Beyond the figure-eight and 2-sphere:
    • Torus: $\pi_{1}(T^{2})\simeq\mathbb{Z}\times\mathbb{Z}$ via encode-decode
    • Klein bottle: $\pi_{1}(K)\simeq\mathbb{Z}\rtimes\mathbb{Z}$ (semidirect product)
    • Real projective plane: $\pi_{1}(\mathbb{RP}^{2})\simeq\mathbb{Z}_{2}$
    • Orientable surfaces: $\pi_{1}(\Sigma_{g})\simeq\langle a_{1},b_{1},\ldots,a_{g},b_{g}\mid\prod[a_{i},b_{i}]=1\rangle$
    • Möbius band and cylinder: $\pi_{1}\simeq\mathbb{Z}$ via deformation retract to circle
  7. (7)
    Fibrations and fiber sequences: Full fibration theory for computational paths:
    • Fiber types: $\mathrm{Fiber}(f,b):=\{a:A\mid f(a)=b\}$ with projection and inclusion
    • Fiber sequences $F\to E\to B$ with equivalence to fibers
    • Type families as fibrations with path lifting via transport
    • Connecting map $\partial:\pi_{1}(B,b)\to F$ via transport
    • Long exact sequence: $\pi_{1}(F)\to\pi_{1}(E)\to\pi_{1}(B)\to\pi_{0}(F)$
    • Induced maps $f_{*}:\pi_{1}(A,a)\to\pi_{1}(B,f(a))$ with functoriality
    • Exactness proofs using $\Sigma$-type path characterization
  8. (8)
    Suspension-loop adjunction: Foundations for the adjunction $\Sigma X,Y]_{*}\cong[X,\Omega Y]_{*}$:
    • Pointed types and pointed maps with basepoint preservation
    • Suspension $\Sigma X$ as pointed type (north as basepoint)
    • Loop space $\Omega Y$ as pointed type (refl as basepoint)
    • Adjunction map: $f:\Sigma X\to Y$ yields $X\to\Omega Y$ via meridians
    • Basepoint preservation: $x_{0}\mapsto\rho$ when $f(\mathrm{north})=y_{0}$
    • Connectivity structures: IsPathConnectedPointed, IsSimplyConnected
  9. (9)
    Eilenberg-MacLane spaces $K(G,n)$: Characterization of classifying spaces:
    • Group and abelian group structures (GroupStr, AbelianGroupStr)
    • $K(G,1)$ characterization via IsKG1: path-connected, $\pi_{1}\cong G$, $\pi_{2}=1$
    • Circle is $K(\mathbb{Z},1)$: uses encode-decode with $\pi_{2}$ triviality from contractibility3
    • Loop space property: $\Omega(K(G,n+1))\simeq K(G,n)$ (stated)
    • Classifying spaces: $BG=K(G,1)$ structure
8.4. Remaining Future Work
  1. (1)
    Cubical type theory comparison: Relate computational paths to cubical approaches [1], exploring whether explicit rewrites complement cubical computation.
  2. (2)
    Fully constructive encode: Implement the code family to eliminate encode axioms, via cubical embedding or macro-generated eliminators.
  3. (3)
    Syntactic paths: Implement paths as syntactic terms (rather than equality-proof wrappers) to recover non-trivial $\sim$ distinctions.
  4. (4)
    Covering space classification: Prove the Galois correspondence between connected covers and subgroups of $\pi_{1}$.
  5. (5)
    Hopf fibration completion: The Hopf fibration framework is established with $\pi_{1}(S^{3})=1$ proved via SVK. Remaining work includes: constructing the Hopf map explicitly via the join $S^{1}*S^{1}\simeq S^{3}$; proving the equivalence $S^{3}\simeq\Sigma_{x:S^{2}}S^{1}$ without axioms; and extending to the quaternionic ($S^{3}\to S^{7}\to S^{4}$) and octonionic ($S^{7}\to S^{15}\to S^{8}$) Hopf fibrations.